OpenStudy (anonymous):

f(x)= 2x-2, graph just f

OpenStudy (anonymous):

\[y=2x-2\] is a line with slope \(2\) and \(y\) intercept \(-2\)

OpenStudy (anonymous):

What is an inverse function ?

OpenStudy (anonymous):

this a general question?

OpenStudy (anonymous):

given a one to one function \(f\) then the inverse \(f^{-1}\) is the function for which \[f\circ f^{-1}(x)=f^{-1}\circ f(x)=x\] put another way, if \(f(a)=b\) then \(f^{-1}(b)=a\)

OpenStudy (anonymous):

No, I would like to know so I can understand my questions for math better

OpenStudy (anonymous):

lets use the example of your function \[f(x)=2x-2\] then since \[f(3)=2\times 3-2=4\] we would have \[f^{-1}(4)=3\]

OpenStudy (anonymous):

and in fact we could find the inverse easily if you start with \[y=2x-2\] and switch \(x\) and \(y\) , which is what the inverse does, you would get \[x=2y-2\] solving for \(y\) igve s \[x=2y-2\\x+2=2y\\ y=\frac{x+2}{2}=\frac{1}{2}x+1\]

OpenStudy (anonymous):

Where did u get the three for f(3)?

OpenStudy (anonymous):

i made it up

OpenStudy (anonymous):

it was an example in other words